OpenStudy (anonymous):

3.7 = 3 + cos (thetha - 30 degrees) find the value of thetha

OpenStudy (phi):

use x for \(\theta\), it is easier to type 3+cos(x-30º)= 3.7 subtract 3 from both sides 3-3 + cos(x-30º)= 3.7-3 cos(x-30º)= 0.7 do you have a calculator to find the inverse cosine of 0.7?

OpenStudy (anonymous):

yeah but what about the 30 degrees/?

OpenStudy (phi):

say you had the problem cos(y) = 0.7 what is y?

OpenStudy (phi):

once you find y we can say x-30 = y this means x= y+30

OpenStudy (anonymous):

so would the answer be 45.57 degrees

OpenStudy (phi):

acos(y)= 0.7 y= 45.57 (also y= 314.427..) now you must find x use x-30= y or x= 30+y

OpenStudy (anonymous):

45.57+30 thanks